Operations Report – January 9th

Crew 290 Operations Report 09-01-2023

SOL: 2

Name of person filing report: Anna Tretiakova

Non-nominal systems: Suit 8

Notes on non-nominal systems: More details in the “Summary of Suits and Radios” section of the report.


Spirit rover used: Yes

Hours: 241.0

Beginning Charge: 100

Ending Charge: 76%

Currently Charging: yes

Opportunity rover used: Yes

Hours: 154.5

Beginning Charge: 100

Ending Charge: 70%

Currently Charging: yes

Curiosity rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Perseverance rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

General Notes on Rovers: Rear view mirrors on spirit and opportunity is very wobbly and violently shakes when the rover is driven.

Summary of Hab Operations:

WATER USE: 4.9 gallons

Water (static tank): 419.5 gallons

Static tank pipe heater (on or off): on

Static tank heater (on or off): on

Toilet tank emptied: no

Summary of Internet: internet is nominal during times 18:00 to 22:00

Summary of Suits and Radios: Radios nominal. All suits except suit 8 are nominal. The battery of suit 8 is 13.7 volts, but there seems to be a problem with the wiring. It is kinda different from the other suits but I cannot say for sure, because I did not want to take it apart completely. Is it possible to get a wiring diagram? It is also a possibility that one of the components is broken. I am unclear on one of the parts: the small black trapezoid with a lid. It is attached to the positive terminal of the battery.

Summary of Greenhab Operations:

WATER USE: 7 gallons 10oz for Greenhab and 2 gallons for the science dome.

Heater: on

Supplemental Light: off

Harvest: 23g of Dill

Summary of Science Dome Operations:

The pH sensor battle continued: the pregnancy test pH sensor had dead batteries, which were replaced but the sensor still could not calibrate. The pH sensor Thermo Scientific Orion Star A111 malfunction was determined to be a broken wire connection, which was attempted to be fixed unsuccessfully. Tonight during wifi hours, Dr. Tanyeri (crew 290 faculty advisor) will be consulted on this matter. The salinity sensor was tested and it is fully functioning. The salinity of the tap water is 0.4 and the salinity of the Martian soil is 10. This inspires to test the soil in the Greenhab for salinity.

Dual Split Heater or AC: On

Summary of RAM Operations: (Tools used, work done):

Multimeter was used to troubleshoot suit 8. Measuring tape was used to measure the water level. RAM was used as the airlock for EVA 3. Pliers and tweezers were used to attempt repair procedures on the pH sensor Orion Star A111.

Summary of any observatory issues: Nominal

Summary of health and safety issues:

All crew in good health and happy spirits. No safety issues. The kitchen stone and counter was cleaned with soap. The trash in the bathroom was emptied, marked with a “B” and put in the back door airlock. The toilet and sink were then thoroughly cleaned with the cleaning supplies available.

Health and safety note: Black lentils were moldy. This product was not brought in by the MDRS or Crew 290. The package was disposed of in the “NB” trash can. The package was sealed and no other food was contaminated.

Questions, concerns and requests to Mission Support:

Should crew engineer do an EVA repair mission to the rovers and tighten the screws on the rear view mirrors?

What are the battery disposal procedures at MDRS?

May we please have the wiring diagram for the suits? (suit 8 specifically)

Any updates on Shannon’s pH sensor electrodes from the rock shop?

Copyright © The Mars Society. All rights reserved. | Main Site